OpenStudy (anonymous):

if v = 2Mcis pi/4 and w = 10Mcis 2pi/3, where M is a constant, write w/v in polar form

OpenStudy (anonymous):

\[ w = r_1cis(\theta) = r_1e^{i\theta} \\ \ \\ v=r_2cis(\varphi)=r_2e^{i\varphi} \\ \ \\ \ \\ \frac{w}{v} = \frac{ r_1e^{i\theta}}{r_2e^{i\varphi}}=\frac{r_1}{r_2}e^{i(\theta-\varphi)}\]